# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v7

@@ 4.2 Theorem of Ethical Safety Assurance
Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence [29, 30, 31, 32], and to date, there is still no universally accepted and practically actionable rigorous definition [33, 34, 35]. Based on the MLT-MF framework, we are able to theoretically ensure the ethical safety of artificial intelligence through the definition of ethical safety and the corresponding formal expressions. To achieve this, it is first necessary to specify appropriate well-formed formulas $Ec$ as ethical constraints—such as those governing fairness and privacy protection—thus establishing a formalized definition of ethical safety.
@@
Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, $Iq$ and $Ir$ respectively, the MLT-MF framework provides:
@@
Let $S_{or}^{Ec}(or,Tr_{h})=G_{max}$ and $S_{or}^{\sim Ec}(or,Tr_{h})=S_{or}(or,Tr_{h})\backslash S_{or}^{Ec}(or,Tr_{h})$. In addition, introduce a new constant in $\mathcal{L}$: NEc, representing a non-Ec ethical safety prompt, and a new predicate: IsSafeback1(NEc), indicating that M outputs the non-Ec ethical safety prompt NEc. We may safely assume that $\mathcal{L}$ originally contains neither.
@@ 4.3 Theorem on Model Generalization Error Estimation
Improving model generalization ability is among the most critical objectives in machine learning. However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance [40, 41, 42, 43, 44]. It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples. By applying the MLT-MF, it is possible to estimate an upper bound on the model’s generalization error.
Theorem 5 (Upper Bound for Model Generalization Error): Let a model M, after learning and optimization, carry the information $Iu=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$, and let the input information be $Iq=(oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq})$. Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively, and let $p_{q}(x)$ be the probability distribution over $S_{oq}$. The total probability of $p_{q}(x)$ on $S_{ou}\cap S_{oq}$ is denoted as $P_{Mq}$. If $S_{ou}$ and $S_{oq}$ are governed by probability distributions defined on the same measurable space, and M is sufficiently trained on $S_{oq}$ and predicts unknown states using a conservative uniform distribution, then the upper bound of the generalization error when processing $Iq$ is determined by the overlap between $S_{oq}$ and $S_{ou}$, as well as the entropy of $p_{q}(x)$ on the non-overlapping region, specifically:
$\mathcal{E}_{gen}\begin{cases}\to 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.13)
where $\mathcal{E}_{gen}$ denotes the upper bound of generalization error, and
$H(p_{q})=-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)$
is the entropy of the data in the non-overlapping region.
Proof: Let $p_{M}(x)$ be the probability distribution determined by the model structure and parameter set of M. $Iu=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$ is the set of information carried by M, so $p_{M}(x)>0$ for $x\in S_{ou}$; $p_{q}(x)>0$ for $x\in S_{oq}$.
Since M is sufficiently trained on $S_{oq}$ and adopts a conservative uniform distribution for unknown states, $p_{M}(x)$ in the overlapping region $S_{ou}\cap S_{oq}$ approximates the true distribution $p_{q}(x)$, and in the non-overlapping region $S_{oq}\backslash S_{ou}$, it follows a uniform distribution. That is,
$p_{M}(x)\begin{cases}\to p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.14)
Using KL divergence to measure the discrepancy between the true and model distributions [45]:
@@
In the non-overlapping region, since the true distribution $p_{q}(x)$ is arbitrary, again by (4.14):
@@
Because $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can estimate the generalization error of model M using the total variation distance (TVD) and KL divergence. According to Pinsker’s inequality [46]:
$\mathcal{E}_{gen}\leq\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
@@
Thus, the latter part of (4.13) is proved. Since we have assumed that $Iu=(ou,Tu_{h},S_{ou},M,$ $Tu_{m},S_{M})$ and $Iq=(oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq})$ are both recoverable information, and the states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ are the physical realizations of $S_{ou}$ and $S_{oq}$ on the model M and its interface, equation (4.13) indeed characterizes the upper bound of model M’s generalization error. Q.E.D.
@@
Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 5—based on MLT-MF—explicitly quantifies the generalization error upper bound using $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, which reflect the overlap between training data and model information, and the entropy $H(p_{q})$ of the non-overlapping region, thereby offering a new perspective and practical guidance for data collection and model design. In particular, Theorem 5 reveals that as long as the problem space is contained within the model’s information space, the model’s generalization error can be made arbitrarily small in the statistical sense. This leads to a critical principle in machine learning:
@@
Corollary 2 (Generalization Error of Ideal Models): If the model information contains the problem information, then the generalization error for these problems is almost zero.

### v9

@@ 4.2 Theorem of Ethical Safety Assurance
Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence mathsea2002 ; Doshi2017 ; Rubin2019 ; Miller2019 , and to date, there is still no universally accepted and practically actionable rigorous definition Russell2015 ; Jobin2019 ; Corrêa2023 . Based on the MLT-MF framework, we are able to theoretically ensure the ethical safety of artificial intelligence through the definition of ethical safety and the corresponding formal expressions. To achieve this, it is first necessary to specify appropriate well-formed formulas $Ec$ as ethical constraints—such as those governing fairness and privacy protection—thus establishing a formalized definition of ethical safety.
@@
Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, IqIq and $Ir$ respectively, the MLT-MF framework provides:
@@
[…]
@@ 4.3 Theorem on Model Generalization Error Estimation
Improving model generalization ability is among the most critical objectives in machine learning. However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance Vapnik1999 ; Bartlett2002 ; Zhang2016 ; Advani2020 ; Perlaza2024 . It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples.Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
Theorem 6 (Upper Bound on TVD) Let the information describing the model M after learning and optimization be $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and the input information be $I_{q}=\langle oq,Tq_{h},S_{oq},eq,Tq_{m},S_{cq}\rangle$. Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space. Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively. The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$, the distribution over $S_{oq}$ is $p_{q}(x)$,with $p_{q}(x)>0$ for $x\in S_{oq}$ . The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$. When the model M is sufficiently trained on $S_{ou}$ and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
$TVD(p_{q},p_{M})\begin{cases}\approx 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.11)
For the second case, it is assumed that $S_{oq}\setminus S_{ou}\neq\emptyset$, $1-P_{Mq}>0$, and $|S_{oq}|-|S_{ou}\cap S_{oq}|>0$. Here, $H(p_{q})=-\sum_{x\in S_{oq}\setminus S_{ou}}p_{q}(x)\log p_{q}(x)$ is the entropy of the data in the non-overlapping region.
[…]
$p_{M}(x)\begin{cases}\approx p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.12)
Using KL divergence to measure the discrepancy between the true distribution and the hypothesized model distribution, we have Goodfellow2016 :
@@
This establishes the first case in (4.11).
In the non-overlapping region, the true distribution $p_{q}(x)$ can be arbitrary. Similarly, by (4.12), we have
@@
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance. By Pinsker’s inequality Pinsker1964 , we have:
$\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
@@
Since we have already assumed that the carrier states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ of $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and $I_{q}=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model M and its interface, respectively. It follows that equation (4.11) characterizes the upper bound of the model’s TVD. This completes the proof.
Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 6, based on the MLT-MF framework, provides an explicit quantitative estimate for the model’s TVD. This estimate is grounded in the overlap between the test and training data $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, as well as the entropy $H(p_{q})$ of the test data distribution over the non-overlapping region. These metrics serve as indicators of how much the testing data shares information with the model’s prior knowledge and how unpredictable the unseen portion is.
Because TVD characterizes the alignment between distributions at a fundamental level—independent of any particular model architecture or task—it is generally applicable to the analysis of model training performance. In fact, for classification tasks with 0-1 loss, the generalization error upper bound is equivalent to the TVD Rodríguez2021 . For regression tasks, the upper bound on generalization error is given by the product of TVD and the Lipschitz constant of the loss function Devroye2013 . Thus, TVD is a meaningful and informative measure for estimating model generalization error.
@@
This shows that Theorem 6 offers a new perspective and practical guidance for data collection, model design, and training for construction. In particular, Theorem 6 reveals that as long as the problem space is contained within the model’s information space, the model’s TVD can be made arbitrarily close to zero in a statistical sense. This is a fundamental result in machine learning:
@@
Corollary 2 (TVD for Ideal Models) If the model’s information encompasses all relevant information for the problem, then the TVD for these problems is approximately zero.
5 Experimental Results
Using the feedforward neural network model listed in the appendix, we conducted regression experiments on two bivariate root functions: $y_{1}=\sqrt{4x_{1}^{2}+3x_{2}^{2}}$ and $y_{2}=\sqrt{2x_{1}^{2}+x_{2}^{2}}$. Both the training and test datasets were generated randomly. We evaluated the models’ prediction errors under various conditions, assessed the impact and computational cost of ethical and safety mechanisms, and explored the challenge of providing multiple types of explanations for the same output. These experiments serve to illustrate and validate the definitions and theorems presented in Section 4.
5.1 Prediction Error Experiments
To evaluate the model’s predictive performance in both overlapping and non-overlapping cases between the training and test datasets, we used the following data setup: 10,000 training samples were generated with $x_{1}$ and $x_{2}$ independently and uniformly distributed over $[0.5,1.5]$, and 10,000 test samples were drawnuniformly from within a circle of radius 1.0 centered at (1.5, 1.0). The model’s learning rate was set to $\eta=0.01$.
Figure 2 illustrates the experimental conditions and outcomes:
Figure 2: Model Prediction Error Experiments and Results Under Different Conditions. (a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples. Among these, 4,951 samples overlap with the training set. (b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset. (c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound. (d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.
(a) The overlap rate between the test and training datasets is approximately 50%.
(b) For 10 randomly selected samples from the overlapping region, as the number of training steps increases from 10 to 4000 (i.e., as the model becomes more sufficiently trained), the root mean square error (RMSE) of the model’s predictions steadily approaches zero. This result aligns with the first case of Theorem 6, equation (4.11).
(c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used. For almost all samples (except a few outliers), their RMSE is less than $\text{TVD}=\sqrt{((1-0.4951)\ln((10000-4951)/(1-0.4951))-2.9741)/2}\approx 0.9155$. This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.
(d) For 10 randomly selected test samples from the non-overlapping region, when using the model’s prediction method, the RMSE also decreases with more training steps, reflecting a certain degree of model generalization outside the training data region. However, the rate of error reduction is significantly lower than in the overlapping region, and for some samples, the RMSE remains higher than that of the conservative uniform prediction. In fact, when training is insufficient, the overall RMSE of the model’s predictions can even exceed the TVD. Meanwhile, the RMSE of conservative predictions for these samples is much lower than the TVD, reflecting the conclusion of the second case in Theorem 6, equation (4.11).
In summary, this regression experiment using a simple feedforward neural network model demonstrates and validates the accuracy of the model TVD upper bound estimation based on the MLT-MF framework, and its potential for estimating generalization error. Since the theoretical proof of Theorem 6 is independent of specific models and tasks, it promises to be extensible to more complex scenarios.
5.2 Ethical Safety Assurance Experiment
Given the nature of the bivariate root function regression task, the model’s predicted outputs should never be negative. This requirement is analogous to Definition 10, which ensures that model outputs do not violate the ethical constraints . Therefore, using the same training dataset as in the previous prediction error experiments, we set the learning rate to and trained the model for 165 steps. For testing, we generated a dataset of 10,000 samples with drawn uniformly from within a circle of radius 0.5 centered at (-1, -1). The criterion for judging ethical safety violations was whether the model’s predicted output was negative. This setup was used to test the feasibility of Theorem 5 and to evaluate the additional computational overhead introduced by ethical safety checking.
For each trial, 5,000 test samples were randomly selected using the trial number as a random seed. Two types of tests were conducted in parallel: “Without Ethical Safety Check” and “With Ethical Safety Check.” In the “Without Ethical Safety Check” condition, model predictions were output directly and the total computation time was recorded. In the “With Ethical Safety Check” condition, each prediction was first checked for sign; if any prediction was negative, an ethical error message was output, otherwise, the model’s prediction was output. The total computation time was also recorded in this scenario. In total, 10 trials were conducted, and the results are shown in Table 1.
Table 1: Summary Table of Ethical Safety Experiment Results
Trial Computation Computation Computation Number of Number of Correct Correct
Number Time Without Time With Time Ratio Ethically Unethical Prediction Prompt
Ethical Ethical Safe Warning Rate for Rate for
Safety Check Safety Check Samples Samples Safe Unethical
(seconds) (seconds) Outputs Outputs
1 0.367 0.391 1.063 4517 483 100% 100%
2 0.365 0.378 1.036 4521 479 100% 100%
3 0.352 0.362 1.030 4548 452 100% 100%
4 0.330 0.378 1.146 4516 484 100% 100%
5 0.371 0.383 1.032 4523 477 100% 100%
6 0.327 0.380 1.160 4546 454 100% 100%
7 0.358 0.402 1.123 4545 455 100% 100%
8 0.360 0.410 1.140 4538 462 100% 100%
9 0.361 0.365 1.010 4519 481 100% 100%
10 0.355 0.424 1.192 4533 467 100% 100%
In Table 1, the “Computation Time Ratio” is calculated as the computation time with ethical safety checking divided by the computation time without it. For the simple feedforward neural network model used in these experiments, the additional computational overhead for performing ethical safety checks before model output ranged from 1.0% to 19.2%. As model size increases, this overhead is expected to decrease in proportion; however, as the number of ethical safety constraints grows, the overhead will increase accordingly. This is a key consideration for implementing ethical safety in models. In these 10 trials, more than 19% of samples resulted in unethical predictions. In every trial, both the “Correct Prediction Rate for Safe Outputs” and the “Correct Prompt Rate for Unethical Outputs” were 100%, demonstrating that by incorporating safety checks and informative outputs, we can always guarantee the ethical safety of model predictions while maximizing the number of safe outputs.
In summary, this ethical safety experiment, grounded in Definition 10 and the formal MLT-MF-based framework for ethical safety, illustrates and validates Theorem 5 regarding the feasibility of integrating ethical safety checking mechanisms and achieving the maximal number of safe outputs.
5.3 Interpretability Experiment
For our bivariate root regression task, the output should never be negative. However, in the ethical safety assurance experiment, after 165 training steps, the model produced negative predictions for more than 19% of randomly selected 5,000-sample test batches, repeated over 10 trials. According to Theorem 4, this phenomenon can be explained from multiple perspectives, each based on different information.
5.3.1 Mathematical Operation-Based Explanation
After 165 training steps, the neural network model M has the following weight matrices and bias vectors from the input layer to the hidden layer:
$W_{1}\approx\begin{bmatrix}-0.065&-0.429\\ 0.681&-0.157\\ -0.498&-0.586\end{bmatrix},\quad b_{1}\approx\begin{bmatrix}0.129\\ -0.551\\ 0.350\end{bmatrix}$ (5.1)
The weights and biases from the hidden layer to the output layer are:
$W_{2}\approx\begin{bmatrix}0.145&0.519&-0.423\\ 0.022&0.048&0.043\end{bmatrix},\quad b_{2}\approx\begin{bmatrix}0.613\\ 0.199\end{bmatrix}$ (5.2)
Therefore, for the test sample $X=[-1.3535,-1.3535]^{T}$, the model predicts:
$Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\text{Identity}(W_{2}\text{ReLU}(W_{1}X+b_{1})+b_{2})\approx\begin{bmatrix}-0.040\\ 0.277\end{bmatrix}$ (5.3)
Here, $y_{1}=-0.040<0$, which mathematically explains why model M outputs a negative value for input $X$.
These formulas provide the designer’s mathematical representation of the model’s computational process. They are equivalent to the formal forward propagation rules (A.21) and (A.22) in the appendix. According to the OIT perspective, these equations describe the ontological state of the model, its outputs, and the information they convey. The model, along with its input and output mechanisms, serves as the carrier of this information and realizes the computational process. Therefore, using these formulas to explain why the model produces negative outputs satisfies the requirements of Definition 9 regarding model interpretability and meets the needs of researchers who seek to understand model behavior from a mathematical standpoint.
5.3.2 Visualization-Based Explanation of the Computational Process
Depicting the model’s computational process visually can make explanations of model behavior much more intuitive. To this end, for model M after 165 training steps, we illustrate its complete structure—covering the inputs, internal neural network calculations and data flow, and outputs—in Figure 3.
Figure 3: Schematic diagram of the computational process for model M after 165 training steps, showing its handling of the input sample $X=[-1.3535,-1.3535]^{T}$ and output $Y\approx[-0.040,0.277]^{T}$.
Compared with equations (5.1)(5.3), Figure 3 not only comprehensively reflects the input, weight matrices, bias vectors, neuron activation functions, and the specific data and formulas for outputs, but also highlights the interactions between the input/output devices and the model. The figure offers a more direct view of the internal data processing pipeline, making it easier to examine and verify each step of the computation. Thus, the visualization provides a more accessible and concrete explanation of how model M processes inputs and generates results.
5.3.3 Partition-Based Explanation
Figure 4: Illustration of Prediction Polarity and Impact of Training Steps. (a) Polarity distribution of predicted data around the training set region after model M has been trained for 165 steps. The blue square marks the area where the training dataset is located, centered at (1, 1) with a side length of 1. Green dots indicate samples with all-positive predictions; dark green dots indicate samples with all-positive predictions that also belong to the test set; red dots indicate samples with at least one negative prediction that belong to the test set; orange dots indicate samples with at least one negative prediction that do not belong to the test set. (b) Cumulative effect of up to 400 training steps on the prediction polarity at two points: (1, 1) and (-1.3535, -1.3535). The blue and green lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (1, 1) from the training set; the red and pink lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (-1.3535, -1.3535) from the test set.
It is evident that equations (5.1)(5.3) and Figure3 explain only the result of model M processing the specific sample $X=[-1.3535,-1.3535]^{T}$. They do not account for the phenomenon observed in the ethical safety assurance experiment, where each random selection of 5,000 samples from the test dataset yields over 19% negative predictions. To explain this, a broader statistical analysis across more data samples is needed.
Figure 4(a) provides a visual depiction of the decision boundary after 165 training steps, showing where model M predicts negative outputs. Notably, there is a boundary line approximately passing through points (-2.0, 0.4) and (-0.4, -2.0). For samples on the upper right side of this diagonal, model M consistently predicts non-negative values; for those on the lower left side, the model consistently outputs negative values. The test region in the figure, representing the test dataset, falls precisely on the lower left side of this boundary line, which explains why a random selection of 5,000 samples from the test set will always include a substantial portion that triggers negative predictions. This visual and statistical partitioning thus provides an intuitive explanation for the observed frequency of negative outputs in the ethical safety experiment.
Figure 4(a) presents an overall view of numerous inputs processed and predicted by model M. This visualization also represents an ontological state of M as it handles a wider range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping. Thus, according to Definition 9, Figure 4(a) constitutes a valid explanation for the distribution of prediction polarities by M.
5.3.4 Explanation Based on Cumulative Training Effects
The previous three explanations all focus on the behavior of M after 165 training steps. However, model training is a continuous process of cumulative improvement. Examining only the model’s instantaneous performance at a particular training step cannot fully explain the causes of abnormal outputs. To address this, we further trained M for 400 iterations on the training dataset. Figure 4(b) shows the predictions for both $(1,1)$ and (1.3535,1.3535)(-1.3535,-1.3535) after each training step.
It is clear that even for $(1,1)$ — a central sample from the training set — the predicted $y_{2}$ by M remains negative during roughly the first 50 training steps. Only after more than 50 steps does the prediction move away from negative values. For the test sample (1.3535,1.3535)(-1.3535,-1.3535), it takes more than about 180 training steps before the prediction begins to consistently avoid negative values. This cumulative effect is entirely due to the optimization of M’s parameters as training progresses.
In fact, by applying the forward propagation rules (A.21) and (A.22) from the appendix—substituting the model parameters obtained after each training step—one can, through a recursive process, provide a rigorous and step-by-step explanation for M’s predictions on specific input samples after any given amount of training.